Obtain the directional derivative for a scalar field (x,y,z)=3x^2y-y^3z^2 at the point (1,-2,-1) in the direction i+j+k
1
Expert's answer
2020-03-16T13:04:26-0400

Find the gradient at the point "(1,-2,-1)"


"\frac{\partial\Phi}{\partial x}=6xy=6\cdot(-2)\cdot1=-12"


"\frac{\partial\Phi}{\partial y}=3x^2-3y^2z^2=3\cdot 1^2-3 \cdot(-2)^2 \cdot (-1)^2=-9"


"\frac{\partial\Phi}{\partial z}=-2y^3z=-2\cdot(-2)^3\cdot(-1)=16"


"\nabla \Phi(1,-2,-1)=-12\overrightarrow{i}-9\overrightarrow{j}+16\overrightarrow{k}=(-12,-9,16)"


Let "\overrightarrow{u}=u_1\overrightarrow{i}+u_2\overrightarrow{j}+u_3\overrightarrow{k}" be a unit vector. The directional derivative at "(1,-2,-1)" in the direction of "\overrightarrow{u}" is


"D_u\Phi(1,-2,-1)=\nabla\Phi(1,-2,-1)\cdot \overrightarrow{u}="


"=(-12\overrightarrow{i}-9\overrightarrow{j}+16\overrightarrow{k})(u_1\overrightarrow{i}+u_2\overrightarrow{j}+u_3\overrightarrow{k})="


"=-12u_1-9u_2+16u_3"


To find the directional derivative in the direction of the vector "(1,1,1)" , we need to find a unit vector in the direction of the vector "(1,1,1)" . We simply divide by the magnitude of "(1,1,1)".


"\overrightarrow{u}=\frac{(1,1,1)}{\lvert\lvert(1,1,1)\rvert\rvert}=\frac{(1,1,1)}{\sqrt{1^2+1^2+1^2}}=\frac{(1,1,1)}{\sqrt{3}}=(\frac{1}{\sqrt{3}},\frac{1}{\sqrt{3}},\frac{1}{\sqrt{3}})"


So,


"D_u\Phi(1,-2,-1)=-12u_1-9u_2+16u_3="


"=-12\frac{1}{\sqrt{3}}-9\frac{1}{\sqrt{3}}+16\frac{1}{\sqrt{3}}=-\frac{5}{\sqrt{3}}" Answer